# LookAhead Tuning: Safer Language Models via Partial Answer Previews

### v2

@@ Abstract.
Fine-tuning enables large language models (LLMs) to adapt to specific domains, but often compromises their previously established safety alignment. To mitigate the degradation of model safety during fine-tuning, we introduce LookAhead Tuning, a lightweight and effective data-driven approach that preserves safety during fine-tuning. The method introduces two simple strategies that modify training data by previewing partial answer prefixes, thereby minimizing perturbations to the model’s initial token distributions and maintaining its built-in safety mechanisms. Comprehensive experiments demonstrate that LookAhead Tuning effectively maintains model safety without sacrificing robust performance on downstream tasks. Our findings position LookAhead Tuning as a reliable and efficient solution for the safe and effective adaptation of LLMs.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processing
@@ 1. Introduction
Fine-tuning effectively enhances the capabilities of large language models (LLMs), for example, fine-tuning chain-of-thought (Wei et al., 2022) data can improve reasoning abilities (Zhao et al., 2024; Muennighoff et al., 2025), and adapting to domain-specific task data can substantially boost task performance (Zhao et al., 2023; Zhang et al., 2024). However, these performance gains often come at the cost of safety alignment (Yang et al., 2023a; Lermen et al., 2023; Zhan et al., 2023; Yi et al., 2024; Alber et al., 2025). Even benign training data can induce catastrophic degradation of protective mechanisms during vanilla fine-tuning processes (Qi et al., 2023; Yang et al., 2023b; Pelrine et al., 2023; He et al., 2024). This raises a critical challenge: how can we endow models with new capabilities without sacrificing their safety? Existing work highlights several obstacles to achieving this goal, including the neglect of the impact of benign data on model safety (Huang et al., 2024), limited methodological adaptability (Wei et al., 2024), and high computational demands (Yang et al., 2024). These issues underscore the need for a solution that is both effective and resource-efficient.
Figure 1. Our method maintains safety alignment comparable to the seed model by rejecting harmful instructions, while achieving task performance improvements equivalent to vanilla fine-tuning.
@@
To address these challenges, understanding how to mitigate forgetting of the model’s safety knowledge (Luo et al., 2023; Li et al., 2024) is promising. Numerous studies have demonstrated that the initial tokens of a model’s response are key predictors of output safety (Lin et al., 2023; Xu et al., 2024; Qi et al., 2024). Inspired by this, we intuitively propose to preserve the model’s learned safety capabilities by previewing the answer prefixes. The central idea is modifying the training data to preview the answer prefix tokens, thereby reducing the loss on these tokens and minimizing disturbances to the model’s initial tokens. This ensures that the model’s inherent safety mechanisms remain intact, even as it learns new capabilities.
@@ 2.3. Token-Level Fine-Tuning
To provide finer control over how closely the fine-tuned model follows the pre-trained distribution, we adopt a token-level objective (Qi et al., 2024). Specifically, a per-token coefficient $\beta_{t}\in[0,1]$ determines the degree of alignment at token $t$. $\beta_{t}=0$ corresponds to standard vanilla fine-tuning; and $\beta_{t}=1$ encourages the fine-tuned model to approximate the original pre-trained model’s output distribution at token $t$. The loss function is defined as:
@@ 2.4. Key Factors in Fine-Tuning Safety
Prior work has identified several key factors influencing model safety during fine-tuning. Specifically, (Zhang and Wu, 2024) and (Lin et al., 2024) have demonstrated that fine-tuning and alignment predominantly impact the initial few tokens of model outputs (Ji et al., 2025). Furthermore, (Xu et al., 2024) and (Qi et al., 2024) have explored that controlled perturbations of initial tokens can serve as effective defense strategies, fundamentally leveraging the concept of KL divergence. Formally, the difference in safety between the fine-tuned model $\theta$ and the original base model $\theta_{0}$ is negatively correlated ($\rho<0$) with the KL divergence ($D^{(m)}_{\text{KL}}(\mathcal{P}_{\theta}\|\mathcal{P}_{\theta_{0}})$) between their output probability distributions ($\mathcal{P}$) over the first $m$ tokens. This relationship can be formally denoted as:
@@ 4.1. Setup
We fine-tune LLaMA2-7B-Chat (Touvron et al., 2023) on GSM8K (Cobbe et al., 2021) and SAMSum (Gliwa et al., 2019) datasets for utility and safety evaluation. For our LookAhead Tuning method, we set the number of previewed tokens $m=6$ for the True Answer Preview strategy and use the buffer prefix $P=\textit{``Let's solve this problem.''}$ for the Virtual Answer Preview strategy. We compare our method against baselines including the Seed Model, Vanilla Fine-Tuning (FT), SDFT (Yang et al., 2024), and Constrained SFT (Qi et al., 2024). Utility is measured by Accuracy for GSM8K and ROUGE-1 for SAMSum. Safety is evaluated with the HEx-PHI dataset (Qi et al., 2023) under the protocol of (Qi et al., 2024), where Raw Safe Rate (RSR) denotes the proportion of safe responses to direct harmful queries, and Jailbreak Safe Rate (JSR) denotes the proportion of safe responses under adversarial prefilling attacks. We fine-tune the model for 3 epochs with a learning rate of $2\times 10^{-5}$ and the batch size of 64, using AdamW (betas 0.5, 0.999). Training has been conducted on 4×NVIDIA A100 (80GB) with bf16 precision and DeepSpeed ZeRO Stage 2 for efficiency.
@@ 4.2. Main Results
As shown in Table 1, both strategy variants of LookAhead Tuning (True/Virtual Answer Preview) perform well across safety and utility evaluations compared to baselines. The True Answer Preview Strategy, i.e., LookAhead Tuning (True), achieves the second-best performance across datasets in terms of RSR and Utility by solely modifying the instruction with explicit prefix guidance. The Virtual Answer Preview Strategy, i.e., LookAhead Tuning (Virtual), attains state-of-the-art performance by jointly modifying instruction and answer with synthetic prefixes, ranking first across all metrics. These findings highlight the effectiveness of prefix-based previews for maintaining model safety while enabling task adaptation. In addition to strong performance, LookAhead Tuning remains resource-efficient. As detailed in Table 2, our true and virtual answer preview strategies respectively increase FLOPs by only $2.18\%$ and $3.90\%$ compared to Vanilla Fine-Tuning, demonstrating the resource efficiency of our approach. Overall, the true answer preview strategy is particularly suited for low-resource scenarios where revealing a small number of answer tokens is acceptable, while the virtual answer preview strategy offers stronger safety guarantees without exposing true prefixes, although slightly increasing resource consumption compared to the former strategy variant. Together, these two strategy variants, by jointly modifying instructions and answers with prefixes, provide flexible mechanisms that balance safety preservation and task performance.
@@ 4.3. Further Analysis
Fine-tuning safety is closely related to the KL divergence of the early tokens.
To better understand this relationship, we analyze the per-token KL divergence of the first few tokens in responses to harmful prompts, between fine-tuned models and the original seed model on the harmful HEX-PHI dataset (Qi et al., 2023, 2024). As shown in Figure 3, Our method significantly reduces KL divergence for the first few (2 & 4) tokens compared to vanilla fine-tuning. In particular, KL divergence values for 6 and 8 tokens remain similar to baseline models, indicating that LookAhead Tuning primarily constrains the earliest tokens while allowing later tokens to adapt freely. This distribution confirms that our approach minimizes perturbations to critical prefix tokens, thereby enhancing safety. These results align with the theoretical framework introduced in §2.4, which predicts a negative correlation between safety retention and KL divergence of early tokens. In practice, we observe that preserving low KL divergence on initial tokens ($D^{(m)}_{\text{KL}}(\mathcal{P}_{\theta}\|\mathcal{P}_{\theta_{0}})\downarrow$) directly corresponds to stronger safety alignment ($\text{Safe}(\theta)\uparrow$) without hindering downstream performance.
@@
Figure 5. Effectiveness of different prefixes using the virtual answer preview strategy. For X-axis prefixes, p1: “Let’s solve this problem.”; p2: “Let’s deal with this situation.”; p3: “dadjalsasdfghkjzmnb”.
@@ 5. Conclusion and Future Work
Looking forward, we plan to extend LookAhead Tuning in two directions. First, we aim to explore broader architectures, including multimodal LLMs, and test robustness under diverse adversarial settings. Second, we will investigate automated strategies for optimizing the trade-off between task utility and safety alignment, potentially adapting the preview length or prefix type dynamically. Furthermore, while LookAhead Tuning enhances safety more effectively than vanilla fine-tuning and other baselines, it may still introduce slight shifts in model alignment. Thus, an important future direction is to further stabilize safety preservation while simultaneously boosting task performance. Ultimately, our work underscores the promise of lightweight, data-driven fine-tuning strategies for safer real-world deployment of LLMs.
@@
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62576307), Ningbo Natural Science Foundation (2024J020), the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023) (Grant A-8001996-00-00), the Academic Research Fund Tier 1 (FY2025) (Grant T1 251RES2507), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.

### v4

@@ Abstract.
Fine-tuning enables large language models (LLMs) to adapt to specific domains, but often compromises their previously established safety alignment. To mitigate the degradation of model safety during fine-tuning, we introduce LookAhead Tuning, a lightweight and effective data-driven approach that preserves safety during fine-tuning. The method introduces two simple strategies that modify training data by previewing partial answer prefixes, thereby minimizing perturbations to the model’s initial token distributions and maintaining its built-in safety mechanisms. Comprehensive experiments demonstrate that LookAhead Tuning effectively maintains model safety without sacrificing robust performance on downstream tasks. Our findings position LookAhead Tuning [^1] https://github.com/zjunlp/LookAheadTuning. as a reliable and efficient solution for the safe and effective adaptation of LLMs.
@@
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3773966.3779370isbn: 979-8-4007-2292-9/2026/02ccs: Computing methodologies Natural language generationccs: Security and privacy Usability in security and privacyccs: Computing methodologies Reasoning about belief and knowledgeccs: Computing methodologies Natural language processing
@@ 1. Introduction
Fine-tuning effectively enhances the capabilities of large language models (LLMs), for example, fine-tuning chain-of-thought (DBLP:conf/nips/Wei0SBIXCLZ22) data can improve reasoning abilities (DBLP:journals/corr/abs-2411-14405; muennighoff2025s1), and adapting to domain-specific task data can substantially boost task performance (DBLP:journals/corr/abs-2303-18223; DBLP:conf/acl/Zhang0ZX24). However, these performance gains often come at the cost of safety alignment (yang2023shadow; lermen2023lora; zhan2023removing; yi2024vulnerability; alber2025medical). Even benign training data can induce catastrophic degradation of protective mechanisms during vanilla fine-tuning processes (qi2023fine; shallowalign; pelrine2023exploiting; he2024s). This raises a critical challenge: how can we endow models with new capabilities without sacrificing their safety? Existing work highlights several obstacles to achieving this goal, including the neglect of the impact of benign data on model safety (huang2024vaccine), limited methodological adaptability (wei2024assessing), and high computational demands (DBLP:conf/acl/YangPFWCZL24). These issues underscore the need for a solution that is both effective and resource-efficient.
Figure 1. Our method maintains safety alignment comparable to the seed model, while achieving task performance improvements equivalent to vanilla fine-tuning.
@@
To address these challenges, understanding how to mitigate forgetting of the model’s safety knowledge (DBLP:journals/corr/abs-2308-08747; DBLP:conf/emnlp/Li0FT24) is promising. Numerous studies have demonstrated that the initial tokens of a model’s response are key predictors of output safety (Lin2023TheUS; xu2024safedecoding; qi2024safety). Inspired by this, we intuitively propose to preserve the model’s learned safety capabilities by previewing the answer prefixes. The central idea is modifying the training data to preview the answer prefix tokens, thereby reducing the loss on these tokens and minimizing disturbances to the model’s initial tokens. This ensures that the model’s inherent safety mechanisms remain intact, even as it learns new capabilities.
@@ 2.3. Token-Level Fine-Tuning
To provide finer control over how closely the fine-tuned model follows the pre-trained distribution, we adopt a token-level objective (qi2024safety). Specifically, a per-token coefficient $\beta_{t}\in[0,1]$ determines the degree of alignment at token $t$. $\beta_{t}=0$ corresponds to standard vanilla fine-tuning; and $\beta_{t}=1$ encourages the fine-tuned model to approximate the original pre-trained model’s output distribution at token $t$. The loss function is defined as:
@@ 2.4. Key Factors in Fine-Tuning Safety
Prior work has identified several key factors influencing model safety during fine-tuning. Specifically, (zhang2024dissecting) and (lin2024unlocking) have demonstrated that fine-tuning and alignment predominantly impact the initial few tokens of model outputs (ji2025first). Furthermore, (xu2024safedecoding) and (qi2024safety) have explored that controlled perturbations of initial tokens can serve as effective defense strategies, fundamentally leveraging the concept of KL divergence. Formally, the difference in safety between the fine-tuned model $\theta$ and the original base model $\theta_{0}$ is negatively correlated ($\rho<0$) with the KL divergence ($D^{(m)}_{\text{KL}}(\mathcal{P}_{\theta}\|\mathcal{P}_{\theta_{0}})$) between their output probability distributions ($\mathcal{P}$) over the first $m$ tokens. This relationship can be formally denoted as:
@@ 4.1. Setup
We fine-tune LLaMA2-7B-Chat (Touvron2023Llama2O) on GSM8K (cobbe2021training) and SAMSum (samsum) datasets for utility and safety evaluation. For our LookAhead Tuning method, we set the number of previewed tokens $m=6$ for the True Answer Preview strategy and use the buffer prefix $P=\textit{``Let's solve this problem.''}$ for the Virtual Answer Preview strategy. We compare our method against baselines including the Seed Model, Vanilla Fine-Tuning (FT), SDFT (DBLP:conf/acl/YangPFWCZL24), and Constrained SFT (qi2024safety). Utility is measured by Accuracy for GSM8K and ROUGE-1 for SAMSum. Safety is evaluated with the HEx-PHI dataset (qi2023fine) under the protocol of (qi2024safety), where Raw Safe Rate (RSR) denotes the proportion of safe responses to direct harmful queries, and Jailbreak Safe Rate (JSR) denotes the proportion of safe responses under adversarial prefilling attacks. We fine-tune the model for 3 epochs with a learning rate of $2\times 10^{-5}$ and the batch size of 64, using AdamW (betas 0.5, 0.999). Training has been conducted on 4×NVIDIA A100 (80GB) with bf16 precision and DeepSpeed ZeRO Stage 2 for efficiency.
@@ 4.2. Main Results
As shown in Table 1, both strategy variants of LookAhead Tuning (True/Virtual Answer Preview) perform well across safety and utility evaluations compared to baselines. The True Answer Preview Strategy, i.e., LookAhead Tuning (True), achieves the second-best performance in terms of RSR and Utility by solely modifying the instruction with explicit prefix guidance. The Virtual Answer Preview Strategy, i.e., LookAhead Tuning (Virtual), attains state-of-the-art performance by jointly modifying instruction and answer with synthetic prefixes, ranking first across all metrics. These findings highlight the effectiveness of prefix-based previews for maintaining model safety while enabling task adaptation. In addition to strong performance, LookAhead Tuning remains resource-efficient. As detailed in Table 2, our true and virtual answer preview strategies respectively increase FLOPs by only $2.18\%$ and $3.90\%$ compared to Vanilla Fine-Tuning, demonstrating the resource efficiency of our approach. Overall, the true answer preview strategy is particularly suited for low-resource scenarios where revealing a small number of answer tokens is acceptable, while the virtual answer preview strategy offers stronger safety guarantees without exposing true prefixes, although slightly increasing resource consumption. Together, these two strategy variants, by jointly modifying instructions and answers with prefixes, provide flexible mechanisms that balance safety preservation and task performance.
@@ 4.3. Further Analysis
Fine-tuning safety is closely related to KL divergence of early tokens.
To better understand this relationship, we analyze the per-token KL divergence of first few tokens in responses to harmful prompts, between the original and fine-tuned models on the harmful HEX-PHI dataset (qi2023fine; qi2024safety). As shown in Figure 3, our method significantly reduces KL divergence for first few (2 & 4) tokens compared to vanilla fine-tuning. In particular, KL divergence values for 6 and 8 tokens remain similar to baseline models, indicating that LookAhead Tuning primarily constrains earliest tokens while allowing later tokens to adapt freely. This distribution confirms that our method minimizes perturbations to critical prefix tokens, thereby enhancing safety. These results align with the theoretical framework introduced in §2.4, which predicts a negative correlation between safety retention and KL divergence of early tokens. In practice, we observe that preserving low KL divergence on initial tokens ($D^{(m)}_{\text{KL}}(\mathcal{P}_{\theta}\|\mathcal{P}_{\theta_{0}})\downarrow$) directly corresponds to stronger safety alignment ($\text{Safe}(\theta)\uparrow$) without hindering downstream performance.
@@
Figure 5. Effectiveness of different prefixes using the virtual answer preview strategy. p1: “Let’s solve this problem.”; p2: “Let’s deal with this situation.”; p3: “dadjalsasdfghkjzmnb”.
@@ 5. Conclusion and Future Work
Looking forward, we plan to extend LookAhead Tuning in two directions. First, we aim to explore broader architectures, including multimodal LLMs, and test robustness under diverse adversarial settings. Second, we will investigate automated strategies for optimizing the trade-off between task utility and safety alignment, potentially adapting the preview length or prefix type dynamically.
Acknowledgments
This work was supported by the Information Technology Center and State Key Lab of CAD&CG, Zhejiang University, Ant Group and Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph, Ningbo Natural Science Foundation (2024J020), and the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023) (Grant A-8001996-00-00).